OpenStudy (anonymous):

prove that area of an quadrilateral triangle is √3/4 a2 where a is side of an triang

Parth (parthkohli):

You mean equilateral triangle? You can either use trig ratios or Heron's Formula.

OpenStudy (anonymous):

ya can you solve this question and a2 means a square

OpenStudy (anonymous):

??????????? can anyone here who can solve this ques...........

OpenStudy (anonymous):

\[\sqrt{3}\div4timesa \] and a ka square hai

OpenStudy (dumbcow):

 for a 30:60:90 triangle the side ratios are 1:sqrt3 : 2 so height of triangle is (sqrt3/2)a Area = 1/2 base*height = a/2 * (sqrt3/2)a = (sqrt3/4)a^2